Circular No. 7170

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1999 J4 (LINEAR) An apparently asteroidal object reported by LINEAR on May 15.3 UT

(mag 18.2–18.8) with unusual motion was noted on The NEO Confirmation Page, and it has subsequently been reported to be cometary in appearance by several observers, including now M. Elowitz and F. Shelly from May 17 LINEAR observations. P. Pravec, U. Babiaková, and P. Kušnirák (Ondřejov, 0.65-m f/3.6 reflector + V filter) reported a coma diameter of 0′.2 and a tail 0′.6 long in p.a. 160^{o}, and J. Tichá and M. Tichý (Kleť, 0.57m f/5.2 reflector) noted the object to be slightly diffuse (coma diameter ∼ 7′′), on May 16.9. The discovery observation is given below; additional astrometry (May 15–17) appears on MPEC 1999-K06.

1999 UT  α_{2000}  δ_{2000}  m_{2}
May 15.31791  15^{h}51^{m}26.^{s}96  +45^{o}12′11′′.2  18.3

Preliminary parabolic orbital elements by B. G. Marsden:

T = 1999 Oct. 28.036 TT  ω = 90.^{o}878
Ω = 264.309 }2000.0
q = 3.86165 AU  i = 118.822
1999TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{2}
May 12 15^{h}57.^{m}95 +44^{o}25.^{′}5 3.578 4.132 116.^{o}8 12.^{o}6 18.1
22 15 37.86 +46 33.0 3.598 4.103 113.1 13.1 18.1
June 1 15 16.99 +47 59.8 3.646 4.074 107.9 13.7 18.1
11 14 56.71 +48 46.7 3.716 4.047 101.8 14.2 18.2
21 14 38.24 +48 58.5 3.801 4.022  95.2 14.6 18.2

SUPERNOVA 1999ce IN ANONYMOUS GALAXY M. Papenkova and W. D. Li, University of California at Berkeley, on

behalf of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126), report their discovery of an apparent supernova on an unfiltered image taken on May 16.2 UT (mag ∼ 18.3) with the 0.8-m Katzman Automatic Imaging Telescope (KAIT), the object being confirmed on an earlier image taken on May 10.2 (mag ∼ 18.8). SN 1999ce is located at α = 11^{h}28^{m}51^{s}.57, δ = +57^{o}08′04′′.0 (equinox 2000.0), which is ∼ 1′′.5 southeast of the nucleus of the host galaxy. KAIT images of the same field taken on May 1.3 (limiting mag ∼ 19.0), Apr. 26.4 (limiting mag ∼ 19.0) and Jan. 11.5 (limiting mag ∼ 19.3) show nothing at the position of the new star.

1999 May 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green